OpenStudy (anonymous):

HELPPPPPP !!!!! What is the sum of the interior angles of a nonagon?1800° 1260° 1440° 1620°

jimthompson5910 (jim_thompson5910):

S = 180(n-2) S = 180(9 - 2) ... plug in n = 9 since a nonagon has 9 sides S = ???

OpenStudy (anonymous):

1260 ??

jimthompson5910 (jim_thompson5910):

yep